Lithium-6 / neutron reaction

6Li + n ⟶ ^{4}He + T + n

Ionisation and subsequent neutralisation of hydrogen impurities

Molecular hydrogen atoms become ionised by electron impact

e^{−} + H_{2} ⟶ H_{2}^{+} + 2e^{−}

Further electron collisions leads to dissociation and production of high energy (> 5eV) neutral hydrogen atoms

e^{−} + H_{2}^{+} ⟶ H + H^{+} + e^{−}

Aneutronic Boron-11 / proton reaction

p +^{11} B ⟶ 3 ^{4}He (8.7 MeV )